# Quantum nonlinear effects in the number-conserving analogue gravity of Bose-Einstein condensates

### v2

@@ IV.2 Dynamics of the background field generating curved spacetime
$\begin{split}\partial_{t}\Big{[}-\frac{1}{\lambda}\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}\Big{]}+\nabla\cdot\Big{[}\rho_{0}\nabla\delta\theta_{1}-\frac{{\bm{v}}_{0}}{\lambda}\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}\\ -\frac{\epsilon}{\lambda}(\nabla\delta\theta_{1})\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}\Big{]}\\ +\frac{1}{\lambda}\left.\frac{\partial f_{1}}{\partial\rho_{c}}\right|_{0}\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}-\left.\frac{\partial f_{1}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\\ -\frac{\epsilon}{\lambda}\left.\frac{\partial^{2}f_{1}}{\partial\rho_{c}\partial\theta_{c}}\right|_{0}\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}\delta\theta_{1}=0.\end{split}\vspace{-6em}$ (53)
We have defined the first-order modification to $U$ as
$\lambda=U+\left.\frac{\partial f_{3}}{\partial\rho_{c}}\right|_{0}$ (54)
neglecting a sub-leading correction of $\mathcal{O}(\epsilon^{2})$, which should be compared with its second-order counterpart in Eq. (36).
@@ V.1 Friedmann-Lemaître-Robertson-Walker background
We will now discuss in more detail how the quantum nonlinear effects coming from the backreaction of the quantized field $\hat{\chi}$ change the leading-order Unruh paradigm metric, which governs perturbations to linear order. To exemplify our arguments in the preceding sections we will consider a solution of the GPE for $\phi_{0}$, such that the leading-order $\mathcal{O}(N)$ density $\rho_{0}$ is constant and the condensate is at rest. This produces a simple analogue of the Friedmann-Lemaître-Robertson-Walker (FLRW) metric of cosmology when $c_{s0}=c_{s0}(t)$, cf. Ref. [51], where the time dependence of the sound speed is implemented by using a Feshbach resonance sweep [44] such that $U=U(t)$, with the constant density enforced by a hard-walled trap of infinite height.
@@
$g^{\rm FLRW}_{\mu\nu}=\sqrt{\frac{\rho_{0}}{U(t)}}\begin{bmatrix}-{c_{s0}^{2}(t)}&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix},$ (55)
where the spatial geometry is conformally flat, according to the usual assumption that the (analogue) universe is isotropic and homogeneous at the cosmological scale. In this model, a decreasing $U(t)$ will produce a spacetime corresponding to an expanding universe, and vice versa. The metric of (55) refers to the background-I, which consists of fluid variables $\rho_{0},\theta_{0}$, that is, solutions of the leading-order GPE (5).
@@
In the next subsection, we study the form of the quantum fluctuation renormalized metric $\mathbf{g}_{\mu\nu}$ when the first-order fluctuation field $\delta\theta_{1}$ perceives a Minkowski or FLRW Unruh paradigm metric $g_{\mu\nu}$.
V.2 Renormalized FLRW Unruh metric
@@
$\mathbf{g}_{tt}=g^{\rm FLRW}_{tt}-\sqrt{\frac{\rho_{0}}{U(t)}}\Big{[}\Big{(}\epsilon U(t)\delta\rho_{1}-\epsilon^{2}{\bm{v}}^{2}_{1}\Big{)}\\ +\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}\Big{(}U(t)\rho_{0}+\epsilon U(t)\delta\rho_{1}\Big{)}-\epsilon^{2}U(t)\rho_{0}\frac{\delta\rho^{2}_{1}}{8\rho^{2}_{0}}\Big{]}\\ =-\sqrt{\frac{\rho_{0}}{U(t)}}\Big{[}U(t)\rho_{0}+\frac{3}{2}\epsilon U(t)\delta\rho_{1}+\frac{3}{8}\epsilon^{2}U(t)\frac{\delta\rho^{2}_{1}}{8\rho_{0}}\Big{]},$ (56)
$\mathbf{g}_{ti}$ $=$ $g^{\rm FLRW}_{ti}-\sqrt{\frac{\rho_{0}}{U(t)}}\Big{[}\epsilon{\bm{v}}_{1i}+\epsilon^{2}\frac{\delta\rho_{1}}{2\rho_{0}}{\bm{v}}_{1i}\Big{]},$ (57)
$\mathbf{g}_{ij}$ $=$ $g^{\rm FLRW}_{ij}+\sqrt{\frac{\rho_{0}}{U(t)}}\Big{[}\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}-\epsilon^{2}\frac{\delta\rho_{1}^{2}}{8\rho_{0}^{2}}\Big{]}\delta_{ij}$ (58)
$=$ $\sqrt{\frac{\rho_{0}}{U(t)}}\Big{[}1+\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}-\epsilon^{2}\frac{\delta\rho_{1}^{2}}{8\rho_{0}^{2}}\Big{]}\delta_{ij}.$
@@
$\begin{split}\partial_{t}\Big{[}-\frac{1}{\lambda(t)}\big{(}\partial_{t}\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\big{)}\Big{]}+\nabla\cdot\Big{[}\rho_{0}\nabla\delta\theta_{1}-\frac{\epsilon}{\lambda(t)}(\nabla\delta\theta_{1})\Big{(}\partial_{t}\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}\Big{]}\\ +\frac{1}{\lambda(t)}\left.\frac{\partial f_{1}}{\partial\rho_{c}}\right|_{0}\Big{(}\partial_{t}\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}-\left.\frac{\partial f_{1}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}-\frac{\epsilon}{\lambda(t)}\left.\frac{\partial^{2}f_{1}}{\partial\rho_{c}\partial\theta_{c}}\right|_{0}\Big{(}\partial_{t}\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}\delta\theta_{1}=0.\end{split}$ (59)
Solving this nonlinear partial differential equation for a given $U(t)$, even for the simplest case of instantaneously switching from zero to a finite $U$, where it simplifies considerably, and proving the convergence of the solution is a nontrivial task. We defer this to a future study. Solving similar classes of partial differential equations is performed for example by the Poincaré-Lighthill-Kuo method [50]. We can however conclude already from the level of complexity of Eq. (59) that the renormalized metric will show features not present in the bare FLRW metric.
@@ VI Conclusion
While probing these correcting nonlinear effects on the standard analogue gravity Unruh paradigm will be experimentally challenging, they might in future experiments become testable. For example by increasing, in the highly controlled way feasible in the quantum optical context, the contact coupling $U$ via Feshbach resonances to simulate quantum many-body effects on a FLRW background as discussed above. This, then, takes the detail to which quantum field dynamics in curved spacetime, and thus the background field concept itself, can be tested to another level of accuracy. In experimental terms, this is contingent on the ability to measure higher-order correlation functions than those of the (standard) density-density variety, as experimentally demonstrated, e.g., in Ref. [53].
@@ VII ACKNOWLEDGMENTS
We show here how the quantum backreaction term $\phi_{2}$ affects the evolution of a quantum-backreacted condensate. According to the fluid dynamical description of the BEC, which we have employed here, the primary two equations governing the dynamics are (21) and (22). These two equations in terms of the classical fluid variables are different from the continuity and Euler equations obtained from the GPE (5) due to the presence of $f_{1}$ and $f_{3}$, which encapsulate the backreaction contribution from the depletion cloud. Here, our primary goal is an estimation of the size and development of these contributions in a realistic model of a BEC which accounts for backreaction. We will to this end consider the model studied previously in [36], where an initially noninteracting Bose gas of uniform density at rest, trapped in a hard-walled one-dimensional container (i.e., in a one-dimensional box trap with no lateral movement permitted by the trapping geometry) of total length $\ell$ extending from $x=-\ell/2$ to $x=+\ell/2$, is subjected to an interaction quench at $t=0$. For $t<0$, the gas is described as solution of the GPE $\phi_{0}$ such that $\rho_{0}=|\phi_{0}|^{2}=$ constant and is at rest. In the noninteracting regime ($U=0$), the density of the depleted cloud $\rho_{\hat{\chi}}=0$, and consequently backreaction is also not present, i.e., $\phi_{2}=0$. For $t\geq 0$, the leading-order condensate order parameter $\phi_{0}$ remains unchanged, and quantization of the field $\hat{\chi}$, along with the condensate-correction described by $\phi_{2}$, describes the coupled evolution of the condensate and non-condensate parts fully analytically [38, 37].
@@
Since the analytical solution for $\phi_{c}$ in terms of $\phi_{0}$ and $\phi_{2}$ is known, it is possible to compute $f_{1}(x,t)$ from (60), setting $\phi_{0}=\exp(-i\mu_{0}t)\sqrt{\rho_{0}}$, where $\mu_{0}=U\rho_{0}$ is the (lowest order) chemical potential, and the lowest-order density $\rho_{0}$ is constant, which defines the bulk coherence length scale via $1/\xi_{0}^{2}=U\rho_{0}=c_{s0}^{2}$, so that all quantities are expressed in powers of $\xi_{0}$.

### v3

@@ IV.2 Dynamics of the background field generating curved spacetime
$\begin{split}\partial_{t}\Big{[}-\frac{1}{\bar{\lambda}}\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+f_{3}^{\prime}\delta\theta_{1}\Big{)}\Big{]}+\nabla\cdot\Big{[}\rho_{0}\nabla\delta\theta_{1}-\frac{{\bm{v}}_{0}}{\bar{\lambda}}\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+f_{3}^{\prime}\delta\theta_{1}\Big{)}\\ -\frac{\epsilon}{\bar{\lambda}}(\nabla\delta\theta_{1})\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+f_{3}^{\prime}\delta\theta_{1}\Big{)}\Big{]}\\ +\frac{1}{\bar{\lambda}}\dot{f}_{1}\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+f_{3}^{\prime}\delta\theta_{1}\Big{)}-f_{1}^{\prime}\delta\theta_{1}-\frac{\epsilon}{\bar{\lambda}}\dot{f}^{\prime}_{1}\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+f_{3}^{\prime}\delta\theta_{1}\Big{)}\delta\theta_{1}=0.\end{split}\vspace{-6em}$ (53)
We have defined another modified coupling as follows
$\bar{\lambda}=\lambda+\dot{f}_{3}$ (54)
neglecting a sub-leading correction of $\mathcal{O}(\epsilon^{2})$. This should be compared with its counterpart $\tilde{\lambda}$ in Eq. (36).
@@ V.1 Friedmann-Lemaître-Robertson-Walker background
We will now discuss in more detail how the quantum nonlinear effects coming from the backreaction of the quantized field $\hat{\chi}$ change the leading-order Unruh paradigm metric, which governs perturbations to linear order. To exemplify our arguments in the preceding sections we will consider a solution of the GPE for $\phi_{0}$, such that the leading-order $\mathcal{O}(N)$ density $\rho_{0}$ is constant and the condensate is at rest. This produces a simple analogue of the Friedmann-Lemaître-Robertson-Walker (FLRW) metric of cosmology when $c_{s0}=c_{s0}(t)$, cf. Ref. [51], where the time dependence of the sound speed is implemented by using a Feshbach resonance sweep [44] such that $\lambda=\lambda(t)$, with the constant density enforced by a hard-walled trap of infinite height.
@@
$g^{\rm FLRW}_{\mu\nu}=\sqrt{\frac{\rho_{0}}{\lambda(t)}}\begin{bmatrix}-{c_{s0}^{2}(t)}&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix},$ (55)
which is spacetime-conformally flat, according to the usual assumption that the (analogue) universe is isotropic and homogeneous at the cosmological scale. The purely spatial part of the metric is by itself also flat because the ultracold quantum gas is probed in a nonrelativistic laboratory, i.e. in Euclidean space. In this model, a decreasing $\lambda(t)$ will produce a spacetime corresponding to an expanding universe, and vice versa. The metric of (55) refers to the background-I, which consists of fluid variables $\rho_{0},\theta_{0}$, that is, solutions of the leading-order GPE (5).
@@
In the next subsection, we study the form of the quantum fluctuation renormalized metric $\mbox{\sf g}_{\mu\nu}$ when the first-order fluctuation field $\delta\theta_{1}$ perceives a Minkowski or FLRW Unruh paradigm metric $g_{\mu\nu}$.
V.2 Renormalized FLRW metric
@@
$\mbox{\sf g}_{tt}=g^{\rm FLRW}_{tt}-\sqrt{\frac{\rho_{0}}{\lambda(t)}}\Big{[}\Big{(}\epsilon\lambda(t)\delta\rho_{1}-\epsilon^{2}{\bm{v}}^{2}_{1}\Big{)}\\ +\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}\Big{(}\lambda(t)\rho_{0}+\epsilon\lambda(t)\delta\rho_{1}\Big{)}-\epsilon^{2}\lambda(t)\rho_{0}\frac{\delta\rho^{2}_{1}}{8\rho^{2}_{0}}\Big{]}\\ =-\sqrt{\frac{\rho_{0}}{\lambda(t)}}\Big{[}\lambda(t)\rho_{0}+\frac{3}{2}\epsilon\lambda(t)\delta\rho_{1}+\frac{3}{8}\epsilon^{2}\lambda(t)\frac{\delta\rho^{2}_{1}}{8\rho_{0}}\Big{]},$ (56)
$\mbox{\sf g}_{ti}$ $=$ $g^{\rm FLRW}_{ti}-\sqrt{\frac{\rho_{0}}{\lambda(t)}}\Big{[}\epsilon{\bm{v}}_{1i}+\epsilon^{2}\frac{\delta\rho_{1}}{2\rho_{0}}{\bm{v}}_{1i}\Big{]},$ (57)
$\mbox{\sf g}_{ij}$ $=$ $g^{\rm FLRW}_{ij}+\sqrt{\frac{\rho_{0}}{\lambda(t)}}\Big{[}\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}-\epsilon^{2}\frac{\delta\rho_{1}^{2}}{8\rho_{0}^{2}}\Big{]}\delta_{ij}$ (58)
$=$ $\sqrt{\frac{\rho_{0}}{\lambda(t)}}\Big{[}1+\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}-\epsilon^{2}\frac{\delta\rho_{1}^{2}}{8\rho_{0}^{2}}\Big{]}\delta_{ij}.$
@@
$\begin{split}\partial_{t}\Big{[}-\frac{1}{\bar{\lambda}(t)}\big{(}\partial_{t}\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+f_{3}^{\prime}\delta\theta_{1}\big{)}\Big{]}+\nabla\cdot\Big{[}\rho_{0}\nabla\delta\theta_{1}-\frac{\epsilon}{\bar{\lambda}(t)}(\nabla\delta\theta_{1})\Big{(}\partial_{t}\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+f_{3}^{\prime}\delta\theta_{1}\Big{)}\Big{]}\\ +\frac{1}{\bar{\lambda}(t)}\dot{f}_{1}\Big{(}\partial_{t}\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+f_{3}^{\prime}\delta\theta_{1}\Big{)}-f_{1}^{\prime}\delta\theta_{1}-\frac{\epsilon}{\bar{\lambda}(t)}\dot{f}^{\prime}_{1}\Big{(}\partial_{t}\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+f_{3}^{\prime}\delta\theta_{1}\Big{)}\delta\theta_{1}=0.\end{split}$ (59)
Solving this nonlinear partial differential equation for a given $\lambda(t)$, even for the simplest case of instantaneously switching from zero to a finite $\lambda$, where it simplifies considerably, and proving the convergence of the solution is a nontrivial task. We defer this to a future study. Solving similar classes of partial differential equations is performed for example by the Poincaré-Lighthill-Kuo method [50]. We can however conclude already from the level of complexity of Eq. (59) that the renormalized metric will show features not present in the bare FLRW metric.
@@ VI Conclusion
While probing these correcting nonlinear effects beyond the standard analogue gravity Unruh paradigm will be experimentally challenging, they might in future experiments become testable. For example by increasing, in the highly controlled way feasible in the quantum optical context, the contact coupling $\lambda$ via Feshbach resonances to simulate quantum many-body effects on a FLRW background as discussed above. This, then, takes the detail to which quantum field dynamics in curved spacetime, and thus the background field concept itself, can be tested to another level of accuracy. In experimental terms, this is contingent on the ability to measure higher-order correlation functions than those of the (standard) density-density variety, as experimentally demonstrated, e.g., in Ref. [53].
@@ VII ACKNOWLEDGMENTS
We show here how the quantum backreaction term $\phi_{2}$ affects the evolution of a quantum-backreacted condensate. According to the fluid dynamical description of the BEC, which we have employed here, the primary two equations governing the dynamics are (21) and (22). These two equations in terms of the classical fluid variables are different from the continuity and Euler equations obtained from the GPE (5) due to the presence of $f_{1}$ and $f_{3}$, which encapsulate the backreaction contribution from the depletion cloud. Here, our primary goal is an estimation of the size and development of these contributions in a realistic model of a BEC which accounts for backreaction. We will to this end consider the model studied previously in [36], where an initially noninteracting Bose gas of uniform density at rest, trapped in a hard-walled one-dimensional container (i.e., in a one-dimensional box trap with no lateral movement permitted by the trapping geometry) of total length $\ell$ extending from $x=-\ell/2$ to $x=+\ell/2$, is subjected to an interaction quench at $t=0$. For $t<0$, the gas is described as solution of the GPE $\phi_{0}$ such that $\rho_{0}=|\phi_{0}|^{2}=$ constant and is at rest. In the noninteracting regime ($\lambda=0$), the density of the depleted cloud $\rho_{\hat{\chi}}=0$, and consequently backreaction is also not present, i.e., $\phi_{2}=0$. For $t\geq 0$, the leading-order condensate order parameter $\phi_{0}$ remains unchanged, and quantization of the field $\hat{\chi}$, along with the condensate-correction described by $\phi_{2}$, describes the coupled evolution of the condensate and noncondensate parts analytically [38, 37].
@@
Since the analytical solution for $\phi_{c}$ in terms of $\phi_{0}$ and $\phi_{2}$ is known, it is possible to compute $f_{1}(x,t)$ from (60), setting $\phi_{0}=\exp(-i\mu_{0}t)\sqrt{\rho_{0}}$, where $\mu_{0}=\lambda\rho_{0}$ is the (lowest order) chemical potential, and the lowest-order density $\rho_{0}$ is constant, which defines the bulk coherence length scale via $1/\xi_{0}^{2}=\lambda\rho_{0}=c_{s0}^{2}$, so that all quantities are expressed in powers of $\xi_{0}$.